Can you please help me solve this problems:
1.) Erwin is 5 times older than john and John is twice as old as his sister, Alice. In 2 years time, the sum of their ages will be 58. How old is john now?
2.) John is twice as old as his friend peter. Peter is 5 years older than Alice. In 5 years, John will be three times as old as Alice. How old is Peter?
3.) Reggie is 5 years older than Lani. Five years ago, the sum of their ages is 25. How old are they now?
: Can you please help me solve this problems:
1.) Erwin is 5 times older than john and John is twice as old as his sister, Alice. In 2 years time, the sum of their ages will be 58. How old is john now?
2.) John is twice as old as his friend peter. Peter is 5 years older than Alice. In 5 years, John will be three times as old as Alice. How old is Peter?
3.) Reggie is 5 years older than Lani. Five years ago, the sum of their ages is 25. How old are they now?

Answer by mangopeeler07(442)

1. Erwin is 5 times older than john and John is twice as old as his sister, Alice.

Erwin=5j
John=j; 2a
Alice=a; 1/2j

Use j, since all of the people can be expressed in terms of j.
In 2 years time, the sum of their ages will be 58.
5j+2+j+2+1/2j+2=58

Combine like terms
6.5j+6=58

Subtract 6
6.5j=52

Divide by 6.5
j=8

Erwin=5(8)
John=(8)
Alice=1/2(8)

5(8)+(8)+1/2(8)+6=58
40+8+4+6=58

Erwin=40
*John=8*
Alice=4
----------------------------------------------------------------------------
2. John is twice as old as his friend peter. Peter is 5 years older than Alice.

John=2p
Peter=p; a+5
Alice=p-5

Use p as the variable.
In 5 years, John will be three times as old as Alice.
2p+5=3(p-5+5)

Simplify
2p+5=3p

Subtract 2p from both sides
5=p

John=2(5)
Peter=(5)
Alice=(5)-5

John=10
*Peter=5*
Alice=0
-------------------------------------------------------------------------
3. Reggie is 5 years older than Lani.

Reggie=L+5
Lani=L

Five years ago, the sum of their ages is 25.
(L+5)-5+(L)-5=25

L+L-5=25
2L-5=25
2L=30
L=15

Reggie=(15)+5
Lani=(15)

Reggie=20
Lani=15